in a chemistry class, 15L of a 13% alchohol solution must be mixed with a 30% solution to get a 15% solution. how many liters of the 30% solution are needed?
Answer by mananth(16946)

percent ---------------- quantity
Alcohol type I 13 ---------------- 15 liters
alcohol type II 30 ---------------- x liters
Mixture 15 ---------------- 20 + x liters

13*15+30x=15(20 +x)
195+30x =300+15x
30x-15x=300-195
15 x = 105
/ 15
x= 7 liters of alcohol type II